36-303: Sampling, Surveys and Society Spring 2011

                 TR 03:00PM -- 04:20PM DH A310
              http://www.stat.cmu.edu/~brian/303
                 http://www.cmu.edu/blackboard

Course Information Instructor: Brian Junker, Statistics 132E Baker Hall 268-8874

    brian@stat.cmu.edu

Office Hours:

Immediately after class (or by appointment).

Teaching Assistant: Jionglin Wu, Statistics FMS 328B

(412) 268-1889

                                           jionglin@stat.cmu.edu

Office Hours:

TBA

Statistics Tutoring Room FMS Building (or by appointment).

Overview

Surveys are ubiquitous in modern life. A recent search for “survey” on Google News yielded over 30,000 hits. The first page or so included hits like: “Survey: Texas Voters like Gov. Rick Perry but don’t want him running for president”, “GoIP’s GO800 Completes Survey: 90% Would Rather Use Text Than a Phone Number”, “U.S. gasoline prices rise to $3.08/gallon—survey”, “Survey Finds Religious Influence Waning in America”, “Bullying is rife at British Airways, union survey claims”.

Social and economic policy in the US and elsewhere are informed—and evaluated—by national surveys such as The National Crime Victimization Survey, The National Assessment of Educational Progress, The Current Population Survey, etc. Marketing research regularly uses survey methods; professional journals like Quirks regularly have articles like “Searching for representative Black households”, “How not to market to Asian-American Youth”, “Comparing online and phone research”, “Keep undesirable respondents out of your online survey”. And survey sampling is an indispensable part of economics, political science, and other social sciences.

Yet survey methodology as we know it today was really only developed in the 20^{th} century, and much of the growth in the use of surveys has occurred in just the past 50 years. At the beginning of the 20^{th} century statisticians were debating whether anything less than a complete census—asking every member of a target population—could be reliable, yet today, samples of 500 or 1,000 individuals from populations of millions of people are routinely used. A reliable survey has many interrelated components: identifying a target population, choosing a sampling method, deciding how to collect data from each survey respondent, writing good survey questions, dealing with missing responses, analyzing the data, making a report.

In this course we will look at the role of sampling and sample surveys in the context of U.S. society and its institutions. We will examine the evolution of survey taking in the United States in the context of its economic, social and political uses. This will eventually lead to discussions about the accuracy and relevance of survey responses, especially in light of various kinds of nonsampling error. All students will be required to design, implement, analyze, and report on a sample survey.

Textbooks, Resources and Statistical Methods

The required text (in the bookstore) for the course is:

• Groves, R. M., Fowler, F. J. Jr., Couper, M. P., Lepkowski, J. M., Singer, E. and Tourangeau, R. (2009). Survey Methodology, SECOND EDITION. (paperback). New York: Wiley. ISBN: 978-0-470-46546.

Some other useful resources on survey sampling include (these are not in the bookstore):

• Converse, J. M. & Presser, S. (1986). Survey Questions: Handcrafting the Standardized Questionnaire (paperback). Sage Publications Inc. ISBN: 978-0803927438.

• Kalton, G. (1983). Introduction to Survey Sampling (paperback). Sage Publications, Inc. ISBN: 978-0803921269.

and if you are really interested in survey sampling, the very next serious book you are likely to read is:

• Lohr, S. (2009). Sampling: Design And Analysis. 2^{nd}(hardcover). Duxbury Press. ISBN: 978-0495105275.

There will also be a variety of handouts, examples, and lecture notes at the course website.

Some computation and statistical analysis will be required in the course. You are welcome to use any statistics package you are familiar with, such such as Excel, Minitab, R, Splus, or SAS (all are available in public clusters; for your personal computer, Minitab can be downloaded from www.cmu.edu/myandrew

and R can be downloaded from http://cran.r-project.org/).

You will be using statistical methods you have learned in 36-202 or a similar course. There will be a quiz in the second lecture covering the material you need for this course.

Objectives

1. To gain an appreciation for the uses of sample surveys in American society;

2. To learn the statistical properties associated with sampling from a finite population;

3. To learn about practical aspects of the design and implementation of sample surveys;

4. To understand the nature of non-sampling and measurement errors in surveys and how to control their effects;

5. To design, conduct, analyze, and report on a sample survey.

The course will run in two parallel “tracks”: The “Classwork” track will give you some of the tools you need to do the “Project” track of the course. Both tracks of the class will start right away, to give you enough tools and enough time to successfully complete your projects.

Requirements: Classwork

Classwork will consist of the following four requirements:

• Reading assignments from the textbook and other sources, and taking notes in class;

• Quizzes and classroom discussion;

• Homework assignments (practice with writing, computation, and other tasks);

• Two midterm exams.

Clear, careful writing and interpretation of results is an important part of both weekly homeworks and the projects.

Always be judicious about including computer output and graphs: show enough that we can clearly see what you are doing, but not so much that we will get lost or bored leafing through your work! A good rule of thumb is to remove any figures, tables, graphs, etc. that you do not have something interesting to say about in the text of your homework solution.

Some homework will be submitted to me in class, and some will be submitted electronically by email or on Blackboard. For electronic submissions, only one pdf or text file will be accepted, per assignment. Please do not submit ms word documents; convert them to pdf and submit that instead. Because of font and version differences, ms word may look fine on one computer but be unintelligible on another.

Requirements: Project

For the project you will divide up into groups of about 4–5 students each. Each group will select a survey topic, design questions and a sampling plan, carry out the survey, analyze the results, and produce a final written report.

Clear, careful writing and interpretation of results is an important part of both weekly homeworks and the projects.

Requirements: Grading

The various requirements in the course will be weighted as follows:

Homework, Quizzes, Other Assigments 20%
Midterm (FEB 17) 15%
Final (APR 12) 15%
Citizenship, Discussion, Attendance 10%
Project Phase I  5%
Project Phase II  5%
Project Phase III 30%
Total 100%

Extra Credit

If your group registers for Meeting of the Minds (MoM; Deadline Apr 1) and presents a poster (May 4), I will throw out the lowest non-zero exam score of each person in that group. If your group registers for the Statistics Poster Competition at MoM and wins first place, I will give that group all A’s regardless of their numerical scores. In addition, first place always gets a cash prize from the Statistics Department.

Citizenship

Much of the work done for this course relies on working with others. You are expected to participate fully in both class and in group assignments. You will be graded on this by both the instructor and the other members of your group. Occasionally I will ask you to fill out peer citizenship assessment forms for your group. These evaluations are confidential and anonymous in the sense that only I and the TA will see the results; other students in the class will not. I will, however, provide feedback if necessary. The purpose of the assessment is only to ensure that everyone is participating fully and working together, NOT to assess someones strengths and weaknesses with the class material. Your citizenship grade will be worth 10%.

Plagiarism

Plagiarism is completely unacceptable. For the purposes of this class, plagiarism is using other peoples words or ideas without crediting those people.

• Group work can use the words and ideas generated by that group without giving credit other than listing all group members names on the assignment.

• If you work with someone on an assignment, simply acknowledge that you did so on the first page.

• If you get ideas or words from a website, journal article, etc that you found helpful, cite the source on your assignment/group work, right where you use it. Then put a bibliography or list of sources cited at the end of each piece of work you do.

• You will not be penalized for working together (except for exams) if appropriately cited.

• Each case of plagiarism will be considered individually; consequences will depend on the seriousness of the case.

Communication and Schedule

• \underline{Schedule.}A rough schedule of

– Lecture Topics

– Milestones for the projects

will be distributed in the second or third lecture.

• \underline{Class website.}Please watch the website http://www.stat.cmu.edu/~brian/303 closely. Current lecture notes, assignment deadlines, etc., will be updated there.

• \underline{Blackboard.} Grades will be kept on a Blackboard site for this class.